OpenStudy (anonymous):

find the area of a square in a circle x^2+y^2=9

OpenStudy (anonymous):

radius is 3 ...use πR^2

OpenStudy (anonymous):

x^2+y^2 = R^2

OpenStudy (anonymous):

ok, but what does 9pi equate to? in other words i forgot what piR^2 is a formula for. i've just been using standard form of circle equations....

OpenStudy (accessdenied):

 you could find distance between each point defining the square to be 3sqrt(2) (pythagorean theorem w/ both sides on x- and y- axis being 3 or distance formula), so the area of the square would be A= s^2, s = 3sqrt(2)